# Assortment Optimization for Patient-Provider Matching

### v1

@@ Appendix B Proofs
$\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{*}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]\leq\frac{1}{N}\sum_{i=1}^{N}\pi^{p}(\theta)_{i,j}\theta_{i,j}=p\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\pi^{p}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]$ (25)
See 2
Proof.
We will first prove the reverse direction; that if $G$ is a complete digraph, then $\pi$ and $\pi^{P}$ have the same match rate. Consider patient $i$ in a component that is a connected component of size $k$. This means that $\sum_{j}\pi(\theta)_{i,j}=1$. We will prove the set of providers unmatched and offered to patient $i$ is non-empty; that is $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}>0$ for any ordering $\sigma$ and $\mathbf{y}^{(t)}$ with $\sigma_{t}=i$. $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}=k$ when $\mathbf{y}^{(t)}=\mathbf{1}$, and decreases by one each time a provider from $\pi(\theta)_{\sigma_{t}}$ is selected. For any $j$ with $\pi(\theta)_{i,j}=1$, there exists $k-1$ other $i^{\prime}$ with $\pi(\theta)_{i^{\prime},j}=1$ because all patients within the complete graph have provider $j$ on their assortment. Therefore, at most $k-1$ other patients can select providers from $\pi(\theta)_{i}$, and therefore $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}$ can decrease by at most $k-1$ before $\sigma_{t}=i$. Therefore, $\lVert\pi(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\rVert_{1}\geq k-(k-1)=1>0$. Moreover, because each patient has a non-empty assortment, in expectation, $Np$ patients match, and so both $\pi$ and $\pi^{P}$ achieve $Np$ matches.
Next, we will prove the forward direction: if the match rates are equal between $\pi$ and $\pi^{P}$, then $G$ consists of components that are complete digraphs. We note that $\pi^{P}$ achieves a match rate of $Np$ because each patient is offered a non-empty assortment of size one. Therefore, the only way for $\pi$ to achieve the same match rate is to offer each patient a non-empty assortment. To show this, we first consider some node $u$ in the graph, corresponding to a patient. Suppose that patient $i$ has an assortment of size $k$, indicating that there exist $k$ edges from $i$ to some node. Suppose that there exists a node, $i^{\prime}$ such that $(i^{\prime},i)$ is an edge, but $(i,i^{\prime})$ is not. Then consider the ordering that places $i$ after its neighbors and $i^{\prime}$. In such an ordering, suppose $i^{\prime}$ selects $v_{i}$ and each neighbor of $i$, $u$, select $v_{u}$. This results in neither $i$ nor its neighbors being available when $i$ must select a patient, leaving an empty assortment. Therefore, if patient $i$ has $v_{i^{\prime}}$ on its assortment, then $i^{\prime}$ must have $v_{i}$ on its assortment for the policy $\pi$.
Next, we consider the scenario where there exists a node $w$ such that $w$ is a two-hop neighbor of $i$ but not an immediate neighbor of $i$. Suppose $w$ is adjacent to $i^{\prime}$, so that $i^{\prime}$ is on $w$’s assortment. Order the patients such that $w$ comes first, then $i$’s neighbors, then $i$. Let $w$ select $i^{\prime}$, let $i^{\prime}$ select $i$, and let all of $i$’s other neighbors select themselves. This results in an empty assortment for $i$; therefore, when all assortments are nonempty, there must not exist any two-hop neighbors that are not also one-hop neighbors. In other words, every node in $G$ only has one-hop neighbors. Therefore, in any component, all neighbors are connected, and so $G$ consists of components that are complete digraphs. ∎
@@
Therefore, summing across all providers gives the ratio as
$\frac{1-(1-p)^{N/M}}{2p}$ (28)
See 2
@@
We compute a matrix, $\hat{\mathbf{Y}}$, such that $Y_{i,j}$ corresponds to the probability that patient $i$ is matched with $j$. Then $\mathbb{E}_{\sigma}[\frac{1}{N}\sum_{t=1}^{N}f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}}]=\hat{\mathbf{Y}}\cdot\theta$.
We then need to prove that
$(p*g(f(\mathbf{X}))-\hat{\mathbf{Y}})\cdot\theta\leq 0$ (29)
To prove this, first let $\ell_{u_{i,j}}=(p*g(f(\mathbf{X}))_{i,j}-\hat{Y}_{i,j})$.
We will first bound $\sum_{j=1}^{M}\ell_{u_{i,j}}$ for a fixed $i$. Let $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,2}}\ldots\theta_{i,u_{i,M}}$ for some coefficients $u$. For any $m$, $\sum_{j=1}^{m}\ell_{u_{i,j}}=\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}-\hat{\mathbf{Y}}_{i,u_{i,j}}$. We can next compute $\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}$. Recall from Section 3.4 that we can explicitly compute $g(f(\mathbf{X}))_{i,j}$ through the function $h(n)=\frac{1}{N}\sum_{k=1}^{N}(1-p\frac{n-1}{N-1})^{k-1}$. Then by the of definition $g(f(\mathbf{X})_{i,j})$, we have $p\sum_{j=1}^{m}g(f(\mathbf{X}))_{i,j}=p(\sum_{j=1}^{m}X_{i,u_{i,j}}h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,l}}\rVert_{1})))$.
Next, we note that $\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ represents the probability that any of the top-$m$ options are available and selected. For any provider $j$, the probability that $j$ is available is at least $h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (see Section 3.4). Therefore,
$\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ (30)
@@
$\geq p\sum_{j=1}^{m}X_{i,u_{i,j}}h(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (32)
Combining our statements for $\hat{\mathbf{Y}}$ and $p*g(f(\mathbf{X}))$ gives that $\sum_{j=1}^{m}X_{i,u_{i,j}}\hat{\mathbf{Y}}_{i,u_{i,j}}\geq\sum_{j=1}^{m}h(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})\prod_{l=1}^{m-1}(1-(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})=\sum_{j=1}^{m}p*g(f(\mathbf{X}))_{i,u_{i,j}}$. Therefore, we have that $\sum_{j=1}^{m}\ell_{u_{i,j}}\leq 0$.
Next, we will show that $(p*g(f(\mathbf{X}))_{i}-\hat{\mathbf{Y}}_{i})\cdot\theta_{i}=\sum_{j=1}^{M}\ell_{u_{i,j}}\theta_{i,j}\leq 0$ for all $i$. We will prove this inductively; first note that $l_{i,u_{i,1}}\theta_{i,u_{i,1}}\leq 0$ Next, assume that $\sum_{j=1}^{M-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$, then:
$\sum_{j=1}^{m}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$ (33)
$=\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m}}$ (34)
@@
Therefore, $p*(g(f(\mathbf{X}))_{i}\cdot\theta_{i})\leq 0$ for all $i$, and so summing across all $i$ gives $p*(g(f(\mathbf{X}))\cdot\theta)\leq 0$
See 2
@@
Let $\mathbf{X}=\pi^{P}(\theta)$ for some given $\theta$. Then because $\pi^{P}(\theta)$ is computed by solving a 1-1 bipartite matching problem, we know that $\sum_{i=1}^{N}X_{i,j}\leq 1$ for all $j$ and $\sum_{j=1}^{M}X_{i,j}\leq 1$ for all $i$. Next, we note that $0\leq\lVert\mathbf{X}_{*,j}\rVert\leq 1$, and moreover, if $\mathbf{X}_{i,j}=1$, then $\lVert\mathbf{X}_{*,j}\rVert=1$. Because $h(1)=1$, we have that $f(\mathbf{X})_{i,j}=\mathbf{X}_{i,j}$. Finally, because each row $i$ of $f(\mathbf{X})_{i,j}$ has at most one $j$ with $\mathbf{X}_{i,j}=1$, and because $\bar{\mathbf{X}}_{i,u_{k}}=f(\mathbf{X})_{i,u_{k}}p\prod_{k^{\prime}=1}^{k-1}(1-f(\mathbf{X})_{i,u_{k}})$, we have that $g(f(\mathbf{X}))=f(\mathbf{X})=p\mathbf{X}$. Therefore, $g(f(\mathbf{X}))\cdot\theta$ is exactly p times the value of the bipartite match. Finally,
$\mathbb{E}_{\sigma}[\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\mathbf{X}_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]=p\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}=p(f(g(\mathbf{X}))\cdot\theta).$ (37)
Therefore, the expected match quality for the pairwise policy $\pi^{P}(\theta)$ is the same as $p(\mathbf{X}\cdot\theta)$. ∎
See 4.1
@@
We first note that $G$ is acyclic; if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ and $\theta_{i^{\prime},v_{i^{\prime}}}\leq\theta_{i^{\prime},v_{i}}$, then swapping $v_{i}$ and $v_{i^{\prime}}$ improves the pairwise policy $\pi^{P}(\theta)$. We note that $\pi^{P}(\theta)$ is the optimal solution for $\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)\theta_{i,j}$ by definition, and so such a swap would violate this property, implying that $G$ is acyclic.
@@
    Case 1: There exists no node $w$, between $i^{\prime}$ and $i$ in $\sigma$ so that $w$ and $i^{\prime}$ share an edge (in either direction). Then $i$ and $i^{\prime}$ can be placed in sequence ($i$ comes immideatly before $i^{\prime}$) without impacting the expected reward. This is because all nodes $w$ between $i$ and $i^{\prime}$ have no impact on the preferences of $i^{\prime}$ because $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{w}}$. Moreover, swapping $i$ and $i^{\prime}$ can only increase the expected match quality; placing $i$ before or after $i^{\prime}$ has no impact on the match quality for patient $i^{\prime}$ (as $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{i}}$), but placing $i$ after $i^{\prime}$ can increase the expected match quality if $v_{i}$ goes unmatched.
@@
    Case 2: Suppose there exists a node $w$, such that there is an edge from $w$ to $i^{\prime}$ Then the pair $(w,i^{\prime})$ violates the reverse topological order, and so we can recurse on this. Note that this is a smaller length pair of nodes within the ordering $\sigma^{*}$. Similarly if there is an edge from $i$ to $w$, then $(i,w)$ violates the reverse topological order ,and so we can recurse on this
@@
    Case 4: Suppose that there exists a node $w$ such that there is an edge from $i^{\prime}$ to $w$. There exists no path from $i^{\prime}$ to $i$ due to the acyclic nature of the graph. Additionally, $w$ does not have an edge to $i$ or any of the ancestors of $i$. Move $i^{\prime}$ and all of its descendants before $i$ in the graph; there are no edges from the descendants of $i^{\prime}$ to $i$. Therefore, this flips the order of $i$ and $i^{\prime}$ without impacting any of the descendants; because only $i$ and $i^{\prime}$ are impacted by this change (as the descendants of $i^{\prime}$ are not impacted by $i$), such a change can only increase the expected match quality, as $i$ can potentially match with $v_{i^{\prime}}$.
@@
Next, we will show that the descendants of patient $i$ dictates what patient $i$ selects in $\pi^{A}$. Let $Z_{1},Z_{2},\ldots,Z_{N}$ be a set of Bernoulli random variables, each of which is 1 with probability $p$. If $Z_{t}=1$ then patient $\sigma_{t}$ will select the highest available provider in their assortment, and if $Z_{t}=0$, then patient $\sigma_{t}$ will select no provider. Additionally, let $V_{1},V_{2},\ldots,V_{N}$ be a set of random variables, where $Z_{t}$ denotes the provider selected by patient $\sigma_{t}$ ($V_{t}=0$ if $Z_{t}=0$). Note that $V_{1}$ is only a function of $Z_{1}$; if $Z_{1}=0$, then $V_{1}=0$, and otherwise, $V_{1}=\operatorname*{arg\,max}_{j}\theta_{\sigma_{1},j}$. Next, let $D_{\sigma_{t}}\subseteq[N]$ be the set of descendants for patient $\sigma_{t}$; that is, $d\in D_{\sigma_{t}}$ means there exists a path from $\sigma_{t}$ to $d$ in $G$. Suppose that $V_{\sigma_{t^{\prime}}}=f(\{Z_{d}\}_{d\in D_{\sigma_{t^{\prime}}}})$ for all $t^{\prime}\leq t$. Then patient $\sigma_{t+1}$ will only select some $j$ so $\theta_{\sigma_{t+1},j}\geq\theta_{\sigma_{t+1},v_{\sigma_{t+1}}}$. By the construction of the graph $G$, this corresponds to edges connected to patient $\sigma_{t+1}$, and so $V_{t+1}=g(\{V_{d}\}_{d\in D_{\sigma_{t+1}})}$. Next, note that $V_{d}=f(\{Z_{d^{\prime}}\}_{d^{\prime}\in D_{d}})$. Therefore, we can rewrite $V_{t+1}$ as

### v2

@@ Appendix F Proofs
$p\mathrm{MQ}(\pi^{*}(\theta),\theta,f)\leq\frac{p}{N}\sum_{i=1}^{N}\pi^{P}(\theta)_{i,j}\theta_{i,j}=\mathrm{MQ}(\pi^{P}(\theta),\theta,f)$ (42)
@@
Summing across all providers gives the appropriate ratio. ∎
Theorem 3.7 The following holds for any $\mathbf{X}$ when $f$ is the uniform choice model with probability $p$:
$p\cdot\langle g(h(\mathbf{X})),\theta\rangle\leq\mathrm{MQ}(\mathbf{X},\theta,f)$ (45)
@@
We compute a matrix, $\hat{\mathbf{Y}}$, such that $Y_{i,j}$ corresponds to the probability that patient $i$ is matched with $j$. Then $\mathrm{MQ}(\mathbf{X},\theta,f)=\hat{\mathbf{Y}}\cdot\theta$. We then need to prove that $p\cdot\langle g(h(\mathbf{X})),\theta\rangle-\langle\hat{\mathbf{Y}},\theta\rangle\leq 0$ To prove this, first let $\ell_{u_{i,j}}=(p\cdot g(h(\mathbf{X}))_{i,j}-\hat{Y}_{i,j})$. We will first bound $\sum_{j=1}^{M}\ell_{u_{i,j}}$ for a fixed $i$. Let $\theta_{i,u_{i,1}}\geq\theta_{i,u_{i,2}}\ldots\theta_{i,u_{i,M}}$ for some coefficients $u$. For any $m$, $\sum_{j=1}^{m}\ell_{u_{i,j}}=\sum_{j=1}^{m}p\cdot g(h(\mathbf{X}))_{i,u_{i,j}}-\hat{\mathbf{Y}}_{i,u_{i,j}}$. We can next compute $\sum_{j=1}^{m}p\cdot g(h(\mathbf{X}))_{i,u_{i,j}}$. Recall from Section 3.1.4 that we can explicitly compute $g(h(\mathbf{X}))_{i,j}$ through the function $d(n)=\frac{1}{N}\sum_{k=1}^{N}(1-p\frac{n-1}{N-1})^{k-1}$. Then by the of definition $g(h(\mathbf{X})_{i,j})$, we have $p\sum_{j=1}^{m}g(h(\mathbf{X}))_{i,j}=p(\sum_{j=1}^{m}X_{i,u_{i,j}}d(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,l}}\rVert_{1})))$. Next, we note that $\sum_{j=1}^{m}\hat{\mathbf{Y}}_{i,u_{i,j}}$ represents the probability that any of the top-$m$ options are available and selected. For any provider $j$, the probability that $j$ is available is at least $d(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (see Section 3.1.4). Therefore,
@@
$\geq p\sum_{j=1}^{m}X_{i,u_{i,j}}d(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})\prod_{l=1}^{j-1}(1-(\lVert\mathbf{X}_{*,u_{i,j}}\rVert_{1})$ (47)
Combining our statements for $\hat{\mathbf{Y}}$ and $p\cdot g(h(\mathbf{X}))$ gives that $\sum_{j=1}^{m}X_{i,u_{i,j}}\hat{\mathbf{Y}}_{i,u_{i,j}}\geq\sum_{j=1}^{m}d(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})\prod_{l=1}^{m-1}(1-(\lVert\mathbf{X}_{*,u_{i,m}}\rVert_{1})=\sum_{j=1}^{m}p\cdot g(h(\mathbf{X}))_{i,u_{i,j}}$. Therefore, we have that $\sum_{j=1}^{m}\ell_{u_{i,j}}\leq 0$. Next, we will show that $p\cdot\langle g(h(\mathbf{X}))_{i},\theta_{i}\rangle-\langle\hat{\mathbf{Y}}_{i},\theta_{i}\rangle=\sum_{j=1}^{M}\ell_{u_{i,j}}\theta_{i,j}\leq 0$ for all $i$. We will prove this inductively; first note that $l_{i,u_{i,1}}\theta_{i,u_{i,1}}\leq 0$ Next, assume that $\sum_{j=1}^{M-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}$, then:
$\small\sum_{j=1}^{m}\ell_{u_{i,j}}\theta_{i,u_{i,j}}=\sum_{j=1}^{m-1}\ell_{u_{i,j}}\theta_{i,u_{i,j}}+\ell_{m}\theta_{i,u_{i,m}}$ (48)
@@
Therefore, $p\cdot\langle g(h(\mathbf{X}))_{i},\theta_{i}\rangle\leq 0$ for all $i$, and so summing across all $i$ gives $p\cdot\langle g(h(\mathbf{X})),\theta\rangle\leq 0$
See 3.1
@@
We first prove the match rate equality, then prove the corresponding claim for match quality. For match rate, we note that there are $\sum_{i}\mathbbm{1}[v(\theta)_{i}\geq 0]$ different patients who have a non-empty assortment (that is, patient $i$ has an assortment with $v(\theta)_{i}$). Each patient matches with probability $\frac{\exp(\theta_{i,v(\theta)_{i}})}{\exp(\theta_{i,v(\theta)_{i}})+\exp(\gamma)}$. Summing over all $v(\theta)_{i}\geq 0$, and dividing by $N$ gives the average match rate. We can bound the match quality as
$\leq\max\limits_{\mathbf{X},\sum_{i=1}^{N}X_{i,j}\leq 1,\sum_{j=1}^{M}X_{i,j}\leq 1}\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}X_{i,j}\theta_{i,j}$ (52)
$=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)_{i,j}\theta_{i,j}$ (53)
Next, we note that each edge in the optimal linear program $v(\theta)_{i}$ exists with probability $\frac{\exp(\theta_{i,v(\theta)_{i}})}{\exp(\theta_{i,v(\theta)_{i}}+\exp(\gamma))}$ if $v(\theta)_{i}\geq 0$, and 00 otherwise. Therefore, we get the final claim by summing over all edges and taking the average. ∎
See D.1
@@
We first note that $G$ is acyclic; if $\theta_{i,v(\theta)_{i}}\leq\theta_{i,v_{i^{\prime}}}$ and $\theta_{i^{\prime},v_{i^{\prime}}}\leq\theta_{i^{\prime},v(\theta)_{i}}$, then swapping $v(\theta)_{i}$ and $v_{i^{\prime}}$ improves the pairwise policy $\pi^{P}(\theta)$. We note that $\pi^{P}(\theta)$ is the optimal solution for $\sum_{i=1}^{N}\sum_{j=1}^{M}\pi^{P}(\theta)\theta_{i,j}$ by definition, and so such a swap would violate this property, implying that $G$ is acyclic.
@@
    Case 1: There exists no node $w$, between $i^{\prime}$ and $i$ in $\sigma$ so that $w$ and $i^{\prime}$ share an edge (in either direction). Then $i$ and $i^{\prime}$ can be placed in sequence ($i$ comes immediately before $i^{\prime}$) without impacting the expected reward. This is because all nodes $w$ between $i$ and $i^{\prime}$ have no impact on the preferences of $i^{\prime}$ because $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v_{w}}$. Moreover, swapping $i$ and $i^{\prime}$ can only increase the expected match quality; placing $i$ before or after $i^{\prime}$ has no impact on the match quality for patient $i^{\prime}$ (as $\theta_{i^{\prime},v_{i^{\prime}}}\geq\theta_{i^{\prime},v(\theta)_{i}}$), but placing $i$ after $i^{\prime}$ can increase the expected match quality if $v(\theta)_{i}$ goes unmatched.
@@
    Case 2: Suppose there exists a node $w$, such that there is an edge from $w$ to $i^{\prime}$ Then the pair $(w,i^{\prime})$ violates the reverse topological order, and so we can recurse on this. Note that this is a smaller length pair of nodes within the ordering $\sigma^{*}$. Similarly, if there is an edge from $i$ to $w$, then $(i,w)$ violates the reverse topological order ,and so we can recurse on this
@@
    Case 4: Suppose that there exists a node $w$ such that there is an edge from $i^{\prime}$ to $w$. There exists no path from $i^{\prime}$ to $i$ due to the acyclic nature of the graph. Additionally, $w$ does not have an edge over $i$ or any of the ancestors of $i$. Move $i^{\prime}$ and all of its descendants before $i$ in the graph; there are no edges from the descendants of $i^{\prime}$ to $i$. Therefore, this flips the order of $i$ and $i^{\prime}$ without impacting any of the descendants; because only $i$ and $i^{\prime}$ are impacted by this change (as the descendants of $i^{\prime}$ are not impacted by $i$), such a change can only increase the expected match quality, as $i$ can potentially match with $v_{i^{\prime}}$.
@@
Next, we will show that the descendants of patient $i$ dictate what patient $i$ selects in $\pi^{A}$. Let $Z_{1},Z_{2},\ldots,Z_{N}$ be a set of Bernoulli random variables, each of which is 1 with probability $p$. If $Z_{t}=1$ then patient $\sigma_{t}$ will select the highest available provider in their assortment, and if $Z_{t}=0$, then patient $\sigma_{t}$ will select no provider. Additionally, let $V_{1},V_{2},\ldots,V_{N}$ be a set of random variables, where $Z_{t}$ denotes the provider selected by patient $\sigma_{t}$ ($V_{t}=0$ if $Z_{t}=0$). Note that $V_{1}$ is only a function of $Z_{1}$; if $Z_{1}=0$, then $V_{1}=0$, and otherwise, $V_{1}=\operatorname*{arg\,max}_{j}\theta_{\sigma_{1},j}$. Next, let $D_{\sigma_{t}}\subseteq[N]$ be the set of descendants for patient $\sigma_{t}$; that is, $d\in D_{\sigma_{t}}$ means there exists a path from $\sigma_{t}$ to $d$ in $G$. Suppose that $V_{\sigma_{t^{\prime}}}=f(\{Z_{d}\}_{d\in D_{\sigma_{t^{\prime}}}})$ for all $t^{\prime}\leq t$. Then patient $\sigma_{t+1}$ will only select some $j$ so $\theta_{\sigma_{t+1},j}\geq\theta_{\sigma_{t+1},v_{\sigma_{t+1}}}$. By the construction of the graph $G$, this corresponds to edges connected to patient $\sigma_{t+1}$, and so $V_{t+1}=g(\{V_{d}\}_{d\in D_{\sigma_{t+1}})}$. Next, note that $V_{d}=f(\{Z_{d^{\prime}}\}_{d^{\prime}\in D_{d}})$. Therefore, we can rewrite $V_{t+1}$ as